OpenStudy (anonymous):

What is the slope intercept form that passes through the given point and is perpendicular (-1,-4),9x+3y=8? What is the answer?! Help!! :)

OpenStudy (anonymous):

First we find m for the given equation, which is -3. The second step is to use point-slope: y-y_1=m(x-x_1). M is the slope and x_1 and y_1 are the x and y values of the given point. The equation is then y=(1/3)x-11/3.

OpenStudy (anonymous):

Is that the answer